Find the mean of the following probability of a random variable X. P(1) = 1/8 , P(2) = 3/8, P(4) = 3/8, P(4) = 1/8


1
Expert's answer
2022-03-03T17:23:27-0500

"\mu=1*\frac{1}{8}+2*\frac{3}{8}+3*\frac{3}{8}+4*\frac{1}{8}=\frac{5}{2}=2.5."